a. The orbit of a satellite about the Earth is given by r= 10000 /(1+0.6cos theta)

What is the eccentricity and shape of the orbit? Also calculate the apogee and

perigee distances.


b. A steel ball A collides elastically with another steel ball B (at rest initially) and the

ball B is observed to move off at an angle theta with the initial direction of motion of A.

The mass of B is five times that of A. Determine the direction in which A moves

after collision and the speeds of the two balls.


c. The displacement of an object executing simple harmonic oscillations is given by:

x = 0.02 sin 2pie (t+0.01) m

Determine (a) amplitude of the oscillatory motion, (b) time-period of oscillation,

(c) maximum velocity, (d) maximum acceleration, and (e) initial displacement of

the object.


1
Expert's answer
2021-07-25T09:17:12-0400

Solution.

a. "r=\dfrac{10000}{1+0.6cos\theta}" ;

"r=\dfrac{p}{1+ecos\theta};"

"e=0.6 -" the orbit is elliptical;

"r_p=\dfrac{p}{1+e}; r_a=\dfrac{p}{1-e};"

"r_p=\dfrac{10000}{1+0.6}=6250; r_a=\dfrac{10000}{1-0.6}=25000;"

b. "m_b=5m_a;"

"\theta_b=\theta;"

"m_av_0+0=m_av_acos\theta_a+m_bv_bcos\theta_b;"

"m_av_0=m_av_acos\theta_a+5m_av_bcos\theta;"

"v_o=v_acos\theta_a+5v_bcos\theta;" "cos\theta_a=\dfrac{v_0-5v_bcos\theta}{v_a};"

"\dfrac{m_av_0^2}{2}=\dfrac{m_av_a^2}{2}+\dfrac{5m_av_b^2}{2};"

"v_0^2=v_a^2+5v_b^2;"

c. "x=0.02sin2\pi(t+0.01);"

"x=0.02sin(2\pi t+0.02\pi);"

"A=0.02m;"

"T=\dfrac{2\pi}{2\pi}=1s;"

"v_{max}=\omega A;"

"v_{max}=2\pi\sdot0.02=0.04\sdot3.14=0.1256m/s;"

"a_{max}=\omega^2A;"

"a_{max}=(2\pi)^2\sdot0.02=0.788m/s^2;"

"x_o=0.02sin0.02\pi=0.0012m;"

Answer:"a. e=0.6;r_p=6250; r_a=25000;"

"b." "cos\theta_a=\dfrac{v_0-5v_bcos\theta}{v_a};"

"c. A=0.02m; T=1s; v_{max}=0.1256m\/s; a_{max}=0.788m\/s^2; x_0=0.0012m."